OpenStudy (anonymous):

Find the values of x and y when the similar rectangle show here has the area given. area of smaller rectangle: 20ft squared. Sides of larger rectangle 8ft and 20ft

OpenStudy (anonymous):

Alright, draw both rectangles next to each other and write down the given for both ^_^: We have the following given for the big rectangle which is: L = 20 ft W = 20 ft and the general formula for the area of a rectangle is : \[A = lw = (20)(8) = 160 ft^2\] is it clear now? ^_^

OpenStudy (anonymous):

do you want x and y = sides of the smaler rectangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

smaller*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Starica my life saver yes i want the sides of the smaller rectangle

OpenStudy (anonymous):

Hmm, okay, I think I have found a relationship between the 2 rectangles, I hope this helps. A1 = xy A2 = Lw Since both rectangles are similar then they are proportional , so yu'll have the following relationship: \[A2/A1 = lw/xy = 160/20 = 8\] so : A2 = 8(A1) does this help? :)

OpenStudy (anonymous):

I dont get what your saying. I have to fine what x and y equals. it has to do witch similar figures

OpenStudy (anonymous):

lol, I have showed you the relationship between both rectangles. I'm not sure how to solve it, but I have given you a relationship, can you figure that out from here?

OpenStudy (anonymous):

To find the sides, you have to know the proportionality of the sides and areas

OpenStudy (anonymous):

here what my problems looks like

OpenStudy (anonymous):

Like I have said, since they are similiar then the sides are proportional so you'll have :\[x/y = 8/20\] cross multiply and solve for x or y, I'll solve for x first and you'll get : 8x = 20y x = (20/8 )y x = (5/2)y plug this x in the area of the smaller rectangle: A1 = xy 20 =[ (5/2)y ] y \[20 = (5/2)y^2\] \[y = \sqrt{40/5}=\sqrt{8}= 2\sqrt{2} ft\] now plug this y in the equation we have found which is : \[x = (8/20) y =(8/20)(2\sqrt{2}) = (4\sqrt{2})/5 ft\] Correct me if I'm wrong please ^_^

OpenStudy (anonymous):

I guess I'm wrong since it's not equal to 20 ft^2 :(

OpenStudy (anonymous):

yes gosh i hate geometry

OpenStudy (anonymous):

lol, don't say that, give it a one more try, I'll try again :)

OpenStudy (anonymous):

if you cant im just not going to do it i get a50 or something

OpenStudy (anonymous):

angoo can you help out a bit please? :)

OpenStudy (anonymous):

Lol, I didn't say I can't, I still didn't stop thinkin, I'm going to give it a one more try

OpenStudy (anonymous):

okay, thank you

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

hmm... both rectangles are supposed to be simmilar, right? so the ratio of their width to lenght is the same

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so the first rectangle is 20ftsq = x * z the second is 160ftsq = 8 * 20 so...

OpenStudy (anonymous):

160/20 = x/8 * z/20

OpenStudy (anonymous):

so what does x and y equal of the smaller square

OpenStudy (anonymous):

How is geometry going to help me in my life anyways

OpenStudy (anonymous):

so, we have to find the ratio first, which is, 160/20=8

OpenStudy (anonymous):

life is geometry martin, everything around us is made up of geometric measurements ^_^

OpenStudy (anonymous):

which means that x of the smaller triangle is 8 times smaller than x of the big triangle, and same goes for z

OpenStudy (anonymous):

so..

OpenStudy (anonymous):

x = 8/8 = 1ft z = 20/8 = 2.5ft

OpenStudy (anonymous):

am i wrong mimi?

OpenStudy (anonymous):

hmm, still not equal to 20 ft^2 angoo, but you were close though!

OpenStudy (anonymous):

excellent :)

OpenStudy (anonymous):

WHAT???

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

You got the ratio idea right

OpenStudy (anonymous):

oh, i see now :D

OpenStudy (anonymous):

martin the area of the smaller rectangle is : A = xy and the area is given which is = 20 ft^2 so (x)(y) must be equal = 20.

OpenStudy (anonymous):

yes i know

OpenStudy (anonymous):

And you must know the ratio between both rectangles since both are similar :) which is 8 just like andy has said. Did you get that part?

OpenStudy (anonymous):

so x = 4 and y = 5

OpenStudy (anonymous):

exactly! :)

OpenStudy (anonymous):

since: A2 = 8 A1 then A1 = A2/8 just like what you have done! excellent ^_^

OpenStudy (anonymous):

i am SOOOOO bad at maths :( *crying out loud desperatelly and emmotionally*

OpenStudy (anonymous):

LOL! angoo, you were close ^_^, but martin that doesn't give 20 ft^2 either. X is correct, but what about y?

OpenStudy (anonymous):

LOL :D, just kidding, but my exams to university are approaching, in ONE week

OpenStudy (anonymous):

well thank you guys i really appreciate what you guys have really helped me. I hate high school math and my teacher who cant even teach

OpenStudy (anonymous):

work hard and good luck!

OpenStudy (anonymous):

np :), but did you get the idea martin?

OpenStudy (anonymous):

yes. I hope i get it right its really late and i have a to study for world history and do this homework

OpenStudy (anonymous):

the problem is, I have no idea what to study... there are hundreds of topics :( well, not a big deal anyway, if I fail, I will just join the military...

OpenStudy (anonymous):

study the history then wake up early in the morning when your brain is fresh and you'll figure what (y) is ^_^

OpenStudy (anonymous):

on i cant im not a morning person

OpenStudy (anonymous):

don't say that andy, you must have high hopes and ambitions.

OpenStudy (anonymous):

i have to wake up in 3 hours

OpenStudy (anonymous):

damn it, I would really love to have your optimism

OpenStudy (anonymous):

well you've got to, it has been proven scientifically that the brain mostly absorbs between 4 am - 7 am :) try to wake up during these times :)

OpenStudy (anonymous):

you're a great computer -user, you can do anything!

OpenStudy (anonymous):

just have faith in yourself :)

OpenStudy (anonymous):

and trust your abilities ^_^

OpenStudy (anonymous):

highschool math must of been a breeze for you guys

OpenStudy (anonymous):

well, I could start OpenStudy2, or even FaceBook 2 for that matter :D

OpenStudy (anonymous):

well, I got an A in my final maths exam in high, but forgot half of everything now, just half a year after finishing

OpenStudy (anonymous):

you have no idea martin, I wasn't even that good in math during high school, but I worked hard in my last 2 senior years and have aced the subject ^_^ . I didn't do good because I didn't get it, but when you fully start to understand mathematics, every mathematical problem will seem simple and clear to you :)

OpenStudy (anonymous):

well i hope college isnt this bad, i hope the teachers actually teach

OpenStudy (anonymous):

Don't say that andy, it's there sometwhere in your head but you've got to refresh you mind that's all ^_^, nothing can be forgotton, everything is stored up in the unconscious :)

OpenStudy (anonymous):

they do martin, but try not to outsmart them, they don't like it :)

OpenStudy (anonymous):

well good, cause my teacher cant teach to save her life she ignores our questions

OpenStudy (anonymous):

especially you andy, with your great computer skills, they'll consider you as a threat, just do your work and keep quite ^_^

OpenStudy (anonymous):

I don't trust a WORD that business subject teachers are teaching me

OpenStudy (anonymous):

lol, in school, you can control the teacher :) but in college you can't

OpenStudy (anonymous):

I want to major in some type of computer thing

OpenStudy (anonymous):

AWESOME! :) what are you most into to?

OpenStudy (anonymous):

into*

OpenStudy (anonymous):

because, if they are so good in business, how come they are still teaching and not living in a private island?

OpenStudy (anonymous):

I love ....... :)

OpenStudy (anonymous):

lol, that's the first step in getting a job andy

OpenStudy (anonymous):

im really good at computers like i like to learn about technology and things of that nature

OpenStudy (anonymous):

you love? what? programming ! we know lol

OpenStudy (anonymous):

not exactly.....

OpenStudy (anonymous):

computers and stuff?

OpenStudy (anonymous):

I love something that most people dont understand and think that it is bad and stupid...

OpenStudy (anonymous):

you will martin, set that as your goal and you'll reach it and you'll never fall on your knees in times of hardships, you'll fall standing because you have a dream :)!

OpenStudy (anonymous):

what is it andy? :)

OpenStudy (anonymous):

Be yourself because you were born this way

OpenStudy (anonymous):

I am afraid you will condemn me if I tell you

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

haha (gaga)

OpenStudy (anonymous):

oh come on, you know me, what is it? ^_^

OpenStudy (anonymous):

who cares we are who we are

OpenStudy (anonymous):

I never judge ppl , I accept them as they are :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

ok, well one of my passions is Hacking

OpenStudy (anonymous):

I dont like highschool because people pick on you for the dumbest things

OpenStudy (anonymous):

that's great :)